# Reciprocal Beyond-Diagonal Reconfigurable Intelligent Surface  (BD-RIS):  Scattering Matrix Design via Manifold Optimization

### v3

@@ Abstract
Beyond-diagonal reconfigurable intelligent surfaces (BD-RISs) are emerging as a transformative technology in wireless communications, enabling enhanced performance and quality of service (QoS) of wireless systems in harsh urban environments due to their relatively low cost and advanced signal processing capabilities. Generally, BD-RIS systems are employed to improve robustness, increase achievable rates, and enhance energy efficiency of wireless systems in both direct and indirect ways. The direct way is to produce a favorable propagation environment via the design of optimized scattering matrices, while the indirect way is to reap additional improvements via the design of multiple-input multiple-output (MIMO) beamformers that further exploit the latter “engineered” medium. In this article, the problem of sum-rate maximization via BD-RIS is examined, with a focus on feasibility, namely low-complexity physical implementation, by enforcing reciprocity in the BD-RIS design. We begin by outlining the system model and formulating an optimization problem that aims to enhance the system’s sum-rate by designing a symmetric scattering matrix. In particular, the approach leverages a manifold optimization framework, where a penalty term is added to the objective function to ensure that the symmetry constraint is upheld, with reciprocity further enforced by projecting the obtained solution onto a set of feasible scattering matrices. Simulation results demonstrate the effectiveness of the proposed method in outperforming current state-of-the-art (SotA) approaches in terms of sum-rate maximization.
@@ I Introduction
Intelligent metasurfaces, or reconfigurable intelligent surfaces (RISs), are an emerging resource in the area of wireless communications, offering a new paradigm for controlling foundational features of electromagnetic fields to enhance system performance [1, 2, 3, 4, 5, 6, 7, 8]. These surfaces consist of individual elements, $i.e.$, passive circuits, with controllable properties that allow them to filter and reradiate the incoming signals without amplification, which is accomplished by implementing delays, reducing the amplitude, or changing wave polarization via the circuit’s local reprogrammable impedance.
@@
Thanks to their low-power and reconfigurable nature, the so-called wave-domain signal processing approach offered by RISs offer several key advantages over classical analog/digital signal processing, including ease of deployment and substantial reductions in end-to-end latency and processing delay [9]. Motivated by these advantages, RISs have attracted enormous attention as an effective solution for enhancing wireless channel quality and consequently the communication performance of multiple-input single-output (MISO), multiple-input multiple-output (MIMO), and integrated sensing and communications (ISAC) systems [1]. Indeed, intelligent surfaces revolutionize wireless systems design by allowing dynamic reconfiguring of the propagation environment, achieving substantial gains without the need for additional active radio chains [10, 11, 12, 13].
@@
Early research on the approach predominantly focused on diagonal reconfigurable intelligent surface (D-RIS) architectures, whose scattering matrices enabled simpler analysis and straightforward implementation [14, 15, 16, 17, 18, 19, 20, 21, 22]. Their use has been extensively studied across various communication settings under various design objectives, including sum-rate maximization [14], spectral efficiency improvement [15], energy efficiency enhancement [16], transmit power reduction [17], [18], and physical layer security [19]. To mention a few illustrative examples, in [20] a weighted sum-rate maximization problem under perfect and imperfect channel state information (CSI) was proposed, with algorithms developed for each case. Alternatively, in [21], an active RIS-aided downlink rate-splitting multiple access system is proposed, and its sum-rate performance was analyzed via a novel algorithm. To this extent, a joint optimization problem was formulated which aimed at designing the transmit (TX) beamforming and rate allocation vectors together with the RIS scattering matrix. The approach involves a two-stage design, where fractional programming (FP) techniques were used in combination with quadratic constraint quadratic programming (QCQP) methods to optimize the respective variables. Another excellent example is [19], whereby the authors utilize the D-RIS in an ISAC framework, applying manifold optimization to jointly improve sum-rate performance and enhance security against eavesdropping.
@@
Although the conventional diagonal architecture offers analytical convenience, it also imposes structural constraints that limit the degrees-of-freedom (DoF), restricting its ability to manipulate the signal space. To overcome these limitations, subsequent studies have focused on beyond-diagonal reconfigurable intelligent surface (BD-RIS) architectures, which significantly increase the DoF available for signal shaping [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. Consequently, manifold optimization and fractional programming have been leveraged as powerful tools to improve overall performance gains.
@@
Junior et al. [23] provide a comprehensive tutorial on manifold optimization techniques and their application in RIS-aided massive MIMO (mMIMO) systems. They emphasize that exploiting the geometry of the constraint set typically yields more compelling solutions compared to standard methods of optimization, motivating the approach adopted in this article. Several recent works have applied these principles to specific BD-RIS design problems, employing manifold optimization to enhance secrecy rate [24] and to jointly configure the scattering matrix and TX precoder [25]. Both studies, however, are conducted in a non-reciprocal setting, $i.e.$, without enforcing symmetry.
One strategy for enforcing reciprocity via unitary and symmetry constraints implemented through auxiliary variables and objective penalties is presented in [28]. In that contribution, however, the reciprocity enforcement is somewhat cumbersome, as the constraints are imposed through a decoupling assisted by auxiliary variables, potentially leading to suboptimal results. In contrast, in the approach taken in this article, symmetry is ensured via a direct manifold optimization formulation of the sum-rate maximization problem.
@@
Alternative BD-RIS design strategies extend beyond direct optimization from the sum-rate objective and instead focus on hybrid approaches. For example, in [26] two methods for designing BD-RIS scattering matrices are proposed. The first method seeks to maximize the total signal power received by the users following a passive maximum ratio transmission (MRT) strategy, while the second approach aims at nullifying the multi-user interference via a passive zero forcing (ZF) strategy. In contrast, this work adopts a fundamentally different strategy: rather than decoupling the design into separate MRT or ZF criteria, the scattering matrix is optimized with respect to the sum-rate objective. This direct formulation accounts for both signal enhancement and interference suppression, enabling these aspects to be addressed simultaneously within a unified and coherent optimization framework.
@@
More recently, research has shifted toward non-reciprocal BD-RIS (NRBD-RIS) design within a full-duplex framework [30, 31], with Liu et al. [30] investigating their application in wireless circulator systems. Their results indicate that properly optimized non-reciprocal scattering matrices can yield improvements in both spectral efficiency and interference management compared to existing reciprocal structures, further expanding the applicability of BD-RIS in wireless networks.
Although prior studies, such as [25, 34, 35], incorporate manifold optimization into the scattering matrix design, to the best of our knowledge, no previous work simultaneously addresses the symmetry and unitary constraint of the BD-RIS within a manifold optimization framework directly applied to the sum-rate objective. This novel approach forms the basis for our proposed methodology, which will be discussed in detail in the following sections. All in all, the main contributions of this article are summarized as follows:
  • A novel manifold optimization method with symmetry enforcement is proposed to design the scattering matrix of BD-RIS. The proposed method, dubbed reciprocal BD-RIS (RBD-RIS), is based on a conjugate gradient ascent (CGA) algorithm, and aims to maximize the sum-rate of a BD-RIS-aided multi-user (MU)-MISO system.
  • The closed-form expression for the gradient of the objective function of the aforementioned problem is provided.
  • Thanks to the above, a generalized closed-form solution of the RBD-RIS design problem is given, which can be particularized for group-connected cases of arbitrary sizes, including single- and fully-connected architectures.
  • Simulation results showing that the proposed method outperforms state-of-the-art (SotA) counterparts [26] are provided, which demonstrate that reciprocity in the scattering matrices does not implicate in performance loss when parameterizing BD-RISs.
@@
Organization: Section II introduces the BD-RIS-aided MU-MISO system model, presents the different BD-RIS architectures considered, and formulates the problem considered in this article. Next, a general solution to the first stage of the optimization problem, namely the scattering matrix design is provided in Section III. Simulation results are offerd in Section IV, which help evaluate the performance of the proposed designs. Finally, a brief conclusion is given in Section V, followed by Appendix A and B, where proofs of two Lemmas are detailed.
@@ II System Model
Consider a BD-RIS aided downlink MU-MISO system as illustrated in Figure 1, where a base station (BS) with $N$ TX antennas serves $K$ single receive (RX) antenna users [^1] This scenario is assumed in accordance to SotA counterparts., denoted as $U_{k}$, $\forall k\in\{1,2,\dots,K\}$, and a BD-RIS consisting of $R$ reflective elements (REs). Hence, the BD-RIS scattering matrix is represented by $\mathbf{\Theta}\in\mathbb{C}^{R\times R}$. The channel linking the BS and the BD-RIS ($i.e.$, the BS-BD-RIS channel) is denoted as $\mathbf{H}_{\mathrm{TX}}\in\mathbb{C}^{R\times N}$, and the BD-RIS-$U_{k}$ channel vector is $\mathbf{h}_{k}\in\mathbb{C}^{R\times 1}$.
The transmit signal vector is given by $\mathbf{x}=\mathbf{V}\mathbf{s}$, where the information symbols $\mathbf{s}\in\mathbb{C}^{K\times 1}$ satisfy $\mathbb{E}[\mathbf{s}\mathbf{s}^{H}]=\mathbf{I}$, and the beamforming matrix $\mathbf{V}\in\mathbb{C}^{N\times K}$ meets the power constraint $\|\mathbf{V}\|_{F}^{2}\leq P_{\mathrm{max}}$, with $P_{\mathrm{max}}$ being the maximum transmit power at the BS. It is assumed that the BS-$U_{k}$ channel is blocked [^2] This assumption is made only to enable direct comparison of the proposed technique with SotA methods, $e.g.$ [26], and is not required by our approach, since the scattering matrix design depends only on the cascaded BS-RIS-to-user channels. However, we are aware that overall performance under such assumptions relies on the BS-RIS channel having a sufficiently high rank, which holds only in rich-scattering, large-antenna setups, motivating an extension of the work under poorly scattered channels with LoS links between the BS and users. Such an extension also requires, however, further work on the beamforming algorithm, thus falling beyond the scope of this article, and therefore will be pursued in a follow-up contribution., such that a focus is placed on the performance gains due to the BD-RIS. Thus, the signal $r_{k}\in\mathbb{C}$ received at $U_{k}$ can be expressed as
@@
where $\mathbf{r}\in\mathbb{C}^{K\times 1}$, $\mathbf{H}_{\mathrm{RX}}\in\mathbb{C}^{K\times R}$, and $\mathbf{w}_{n}\in\mathbb{C}^{R\times 1}$, s.t. $n\in\{1,2,\dots,N\}$, denote the received signal vector, BD-RIS-$U_{k}$ channel matrix, and BS-BD-RIS channel vector, respectively.
@@ II-A Scattering Matrix Definition
    Single-connected BD-RIS: Equivalent to the conventional D-RIS, where each RE is independently connected to a single grounded reconfigurable impedance with no interconnections among REs. Consequently, the resulting scattering matrix is inherently diagonal, which naturally satisfies the symmetry constraints, while also being unitary due to the reciprocal and lossless nature. The aforementioned constraints are described as
@@
    Group-connected BD-RIS: In this intermediate architecture the REs are partitioned into $G$ disjoint groups, each consisting of $R_{G}=\frac{R}{G}$ elements. Within each group, all REs are interconnected, enabling waves impinging on one element to be reflected by all others in the same group. Apart from the intra-group connections, no other link exists. Consequently, the scattering matrix can be represented as a block-diagonal matrix by ordering its elements group-by-group, such that each block corresponds to one group and must independently satisfy the unitary and symmetry constraints. Notably, the group-connected architecture generalizes the previous two cases, reducing to the fully-connected case when $G=1$ and to the single-connected case when $G=R$. Denoting the scattering matrix of groups $g\in\{1,2,\ldots,G\}$ by $\mathbf{\Theta}_{g}\in\mathbb{C}^{R_{G}\times R_{G}}$, the constraints on the group-connected BD-RIS scattering matrix can be described as
    $\hskip-21.52771pt\mathcal{S}_{\text{GC}_{1}}\!\!=\!\left\{\mathbf{\Theta}\!:\!\mathbf{\Theta}\!=\!\mathrm{blkdiag}(\mathbf{\Theta}_{1},\ldots,\mathbf{\Theta}_{G}),\!\mathbf{\Theta}_{g}\!=\!\mathbf{\Theta}_{g}^{T},\forall g\right\}\!,\!\!\!$ (7)

### v4

@@ Abstract
Beyond-diagonal reconfigurable intelligent surfaces (BD-RISs) are emerging as a transformative technology in wireless communications, enabling enhanced performance and quality of service (QoS) of wireless systems in harsh urban environments due to their relatively low cost and advanced signal processing capabilities. Generally, BD-RIS systems are employed to improve robustness, increase achievable rates, and enhance energy efficiency of wireless systems in both direct and indirect ways. The direct way is to produce a favorable propagation environment via the design of optimized scattering matrices, while the indirect way is to reap additional improvements via the design of multiple-input multiple-output (MIMO) beamformers that further exploit the latter “engineered” medium. In this article, the problem of sum-rate maximization via BD-RIS is examined, with a focus on feasibility, namely low-complexity physical implementation, by enforcing reciprocity in the BD-RIS design in a manner that adheres to the geometry of the manifold of symmetric matrices. To that end, the sum-rate objective is transformed into a quadratic function via fractional programming (FP), augmented via the also quadratic reciprocity constraint in the form of a regularization term, while the unitary constraint is dealt with via a manifold optimization framework. Simulation results demonstrate the effectiveness of the proposed method in outperforming current state-of-the-art (SotA) approaches in terms of sum-rate maximization.
@@ I Introduction
Intelligent metasurfaces, or reconfigurable intelligent surfaces (RISs), are an emerging resource in the area of wireless communications, offering a new paradigm for controlling foundational features of electromagnetic fields to enhance system performance [2, 3, 4, 5, 6, 7]. These surfaces consist of individual elements, $i.e.$, passive circuits, with controllable properties that allow them to filter and reradiate the incoming signals without amplification, which is accomplished by implementing delays, reducing the amplitude, or changing wave polarization via the circuit’s local reprogrammable impedance.
Thanks to their low-power and reconfigurable nature, the so-called wave-domain signal processing approach offered by RISs offer several key advantages over classical analog/digital signal processing, including ease of deployment and substantial reductions in end-to-end latency and processing delay [8].
@@
Motivated by these advantages, RISs have attracted enormous attention as an effective solution for enhancing wireless channel quality and consequently the communication performance of multiple-input single-output (MISO), multiple-input multiple-output (MIMO), and integrated sensing and communications (ISAC) systems [2]. Indeed, intelligent surfaces revolutionize wireless systems design by allowing dynamic reconfiguring of the propagation environment, achieving substantial gains without the need for additional active radio chains [9, 10].
@@
Early research on the approach predominantly focused on diagonal reconfigurable intelligent surface (D-RIS) architectures, whose scattering matrices enabled simpler analysis and straightforward implementation [11, 12, 13, 14, 15, 16, 17]. Their use has been extensively studied across various communication settings under various design objectives, including sum-rate maximization [11], spectral efficiency improvement [12], energy efficiency enhancement [13], transmit power reduction [14], and physical layer security [15]. To mention a few illustrative examples, in [16] a weighted sum-rate maximization problem under perfect and imperfect channel state information (CSI) was proposed, with algorithms developed for each case. Alternatively, in [17], an active RIS-aided downlink rate-splitting multiple access system is proposed, and its sum-rate performance was analyzed via a novel algorithm. To this extent, a joint optimization problem was formulated to design the transmit (TX) beamforming and rate allocation vectors together with the RIS scattering matrix. The approach involves a two-stage design, where fractional programming (FP) techniques were used in combination with quadratic constraint quadratic programming (QCQP) methods to optimize the respective variables. Another excellent example is [15], whereby the authors utilize the D-RIS in an ISAC framework, applying manifold optimization to jointly improve sum-rate performance and enhance security against eavesdropping.
@@
Although the conventional diagonal architecture offers analytical convenience, it also imposes structural constraints that limit the degrees-of-freedom (DoF), restricting its ability to manipulate the signal space. To overcome these limitations, subsequent studies have focused on beyond-diagonal reconfigurable intelligent surface (BD-RIS) architectures, which significantly increase the DoF available for signal shaping [18, 19, 20, 21, 22, 23, 24, 25, 26]. However, the increase in DoF, achieved through enhanced connectivity of the BD-RIS, is generally accompanied by higher hardware complexity and implementation cost. As a result, a substantial body of literature has investigated the trade-off between achievable performance and hardware complexity in BD-RIS architectures [27, 28].
@@
In this context, advanced optimization techniques have been employed to effectively exploit the additional DoF while respecting structural constraints. Both FP and first-order methods, including projected gradient and accelerated gradient schemes [29], have been widely adopted in RIS optimization problems to improve convergence speed and computational efficiency. While these approaches typically enforce constraints through Euclidean projections onto feasible sets, manifold optimization instead leverages the intrinsic geometry of the constraint space. By operating directly on smooth manifolds via tangent-space projections and retraction steps [30, 31], it maintains feasibility along the optimization trajectory while explicitly accounting for the geometry of the constraint set. Junior et al. [18] provide a comprehensive tutorial on manifold optimization techniques and their application in RIS-aided massive MIMO (mMIMO) systems, where it is shown that adhering to the manifold geometry of the constraints consistently yields better solutions than standard optimization methods, which motivates the approach adopted in this article. Several recent works have applied these principles to specific BD-RIS design problems, employing manifold optimization to enhance secrecy rate [19] and to jointly configure the scattering matrix and TX precoder [20]. Both these as well as similar studies such as [32, 33], however, are conducted for non-reciprocal settings, without enforcing symmetry.
Enforcing a symmetric scattering matrix ensures compatibility with physically realizable passive architectures, avoiding the need for active components, magnetic biasing, or non-reciprocal circuit elements, which increase hardware complexity, power consumption, and implementation cost [7, 24]. One way to enforce the unitary and symmetry constraints inherent to reciprocal BD-RIS is through a penalty dual decomposition (PDD) framework, as proposed in [23], where the scattering matrix and transmit beamformers are jointly optimized for both sum-rate maximization and transmit power minimization problems. While effective, the approach relies on auxiliary-variable decoupling and a double-loop procedure, is primarily tailored to fully- and group-connected BD-RIS architectures, and incurs relatively high computational complexity. To address these limitations, [34] develops an architecture-independent framework based on partially proximal alternating direction method of multipliers (pp-ADMM), which reformulates the BD-RIS constraints via the admittance matrix and enables efficient closed-form or low-complexity updates at each iteration. This general and computationally efficient framework is adopted in this work as a benchmark for performance and complexity comparisons.
@@
Alternative BD-RIS design strategies extend beyond direct optimization from the sum-rate objective and instead focus on hybrid approaches. For example, in [21], two methods for designing BD-RIS scattering matrices are proposed. The first method seeks to maximize the total signal power received by the users following a passive maximum ratio transmission (MRT) strategy, while the second approach aims at nullifying the multi-user interference via a passive zero forcing (ZF) strategy. In contrast, this work adopts a fundamentally different strategy: rather than decoupling the design into separate MRT or ZF criteria, the scattering matrix is optimized with respect to the sum-rate objective. This direct formulation accounts for both signal enhancement and interference suppression, enabling these aspects to be addressed simultaneously within a unified and coherent optimization framework.
@@
More recently, research has shifted toward non-reciprocal BD-RIS (NRBD-RIS) designs, particularly within a full-duplex framework [24, 25]. For example, Liu et al. [24] investigate their application in wireless circulator systems, demonstrating that properly optimized non-reciprocal scattering matrices yield improvements in both spectral efficiency and interference management compared to existing reciprocal structures, further expanding the applicability of BD-RIS in wireless networks. However, the practical realization of non-reciprocal BD-RIS entails increased hardware complexity and resource requirements, which may limit their suitability in large-scale or cost-sensitive deployments. Although non-reciprocal designs can be advantageous in specific scenarios, reciprocal BD-RIS architectures remain highly relevant for the broader range of applications where simplicity, scalability, and energy efficiency are essential.
All in all, we recognize the maximization of sum-rate via the design of reciprocal BD-RIS scattering matrices as a dual manifold problem which, to the best of our knowledge, has not yet been addressed as such via an adequate manifold optimization method, likely due to the challenge of dealing with two manifolds of different geometries simultaneously. Recently, this challenge was addressed in [35], where a Riemannian optimization algorithm operating on the solution space defined by the intersection manifold of unitary and symmetric matrices was proposed. [^1] We emphasize that a pre-print of [35] first appeared only in Jan. 2026 https://arxiv.org/abs/2601.13877, after the submission of this article which was available at http://www.arxiv.org/abs/2509.20246 since Sep. 2025. That formulation is, however, restricted to the fully-connected architecture and primarily considers the single-user case, where inter-user interference is absent.
In contrast to the latter, in this article, the dual-manifold nature of the reciprocal BD-RIS scattering matrix design problem is addressed under a general setting [^2] For the sake of simplicity and to allow direct comparison with state-of-the-art (SotA) methods we explicitly focus on multi-user (MU)-MISO systems, for which closed-form gradients are derived, while emphasizing that the proposed method applies directly to the general MU-MIMO case. See Subsection II-C for more details. by incorporating the reciprocity constraint as a quadratic regulation term into the sum-rate objective transformed also into a quadratic function via FP, while enforcing the unitarity constraint by solving the resulting regularized problem via manifold optimization over the Stiefel manifold. Our main contributions are summarized as follows:
  • Novel technique with improved performance: A novel method is proposed to design the scattering matrix of reciprocal BD-RIS (RBD-RIS), under which both the symmetry and unitary constraints are sensibly enforced, in a manner that accommodates both manifold geometries. In particular, in the proposed method, the quadratic geometry of the symmetry constraint manifold is incorporated into the quadratic sum-rate objective reformulated via FP, while the geometry of the Stiefel unitary constraint manifold is dealt with via manifold optimization. The new method, which has complexity similar to that of SotA techniques, is shown via simulations to outperform the latter under equivalent conditions.
  • Original closed-form expressions: While the method is applicable to general MU-MIMO systems, a closed-form expression for the gradient of the proposed regularized and FP-reformulated objective function is provided for the MU-MISO setup, enabling the implementation of the method via conjugate gradient ascent (CGA) in that case.
  • High hardware compatibility: Thanks to the above, the proposed scattering matrix design of RBD-RIS under MU-MISO settings can be implemented in a highly controlled manner that facilitates efficient hardware design [36]. In particular, with the exception of an Armijo line search, the key steps of the method, namely, the tangential projection, gradient evaluation, and updating of FP variables, are achieved via successive evaluation of closed-form expressions, enabling dedicated signal processing hardware to be optimized. [^3] It is a well-known and widely recognized principle in digital signal processing (DSP) and very large scale integration (VLSI) designs that closed-form algorithms reduce to a fixed sequence of arithmetic operations, favoring the stability, reliability, power consumption and efficacy of hardware implementation, as a result of the corresponding deterministic pipelines, parallelization and static scheduling [36].
@@
Organization: Section II introduces the BD-RIS-aided MU-MISO system model, presents the different architectures considered, and formulates the problem addressed in the article. Next, a general solution to the first stage of the optimization problem, namely the scattering matrix design, is provided in Section III. Simulation results are offered in Section IV, which help evaluate the performance of the proposed designs. Finally, a brief conclusion is given in Section V, followed by Appendix A and B, where proofs of two Lemmas are detailed.
@@ II System Model
Consider a BD-RIS aided downlink MU-MISO system as illustrated in Figure 1, where a base station (BS) with $N$ TX antennas serves $K$ single receive (RX) antenna users [^4] This scenario is considered due to its analytical tractability and simplicity, while enabling direct comparisons with SotA alternatives. The methods can, however, be easily generalized to MU-MIMO settings, as described in Subsection II-C. with the help of a BD-RIS consisting of $R$ reflective elements (REs), The associated BD-RIS scattering matrix is represented by $\mathbf{\Theta}\in\mathbb{C}^{R\times R}$. The channel linking the BS and the BD-RIS ($i.e.$, the BS-to-BD-RIS channel) is denoted as $\mathbf{H}_{\mathrm{TX}}\in\mathbb{C}^{R\times N}$, and the channel vector connecting the BD-RIS and the $k$-th user is $\mathbf{h}_{k}\in\mathbb{C}^{R\times 1}$.
The transmit signal vector is given by $\mathbf{x}=\mathbf{V}\mathbf{s}$, where the information symbols $\mathbf{s}\in\mathbb{C}^{K\times 1}$ satisfy $\mathbb{E}[\mathbf{s}\mathbf{s}^{H}]=\mathbf{I}$, and the beamforming matrix $\mathbf{V}\in\mathbb{C}^{N\times K}$ meets the power constraint $\|\mathbf{V}\|_{F}^{2}\leq P_{\mathrm{max}}$, with $P_{\mathrm{max}}$ being the maximum transmit power at the BS. It is assumed that the channel between the BS and the $k$-th user is blocked [^5] This assumption captures the most challenging and conceptually relevant scenario, where the impact and benefits of RIS-assisted propagation are most clearly observed. In particular, it highlights coverage extension and blockage mitigation, while reducing the need for network-level optimization measures such as deploying additional base stations. As a result, the blocked-user assumption is widely adopted in RIS-related studies, enabling direct comparison of the proposed technique with SotA methods, $e.g.$ [21, 34]. Importantly, this assumption is not required by our proposed approach, since the scattering matrix design depends only on the cascaded BS-to-RIS-to-user channels. Nevertheless, the overall performance under this setting relies on the BS-to-RIS channel having a sufficiently high rank, which holds only in rich-scattering, large-antenna setups, motivating an extension of the work under poorly scattered channels with line-of-sight (LoS) links between the BS and users. Such an extension also requires, however, further work on the beamforming algorithm, thus falling beyond the scope of this article, and therefore will be pursued in a follow-up contribution., such that a focus is placed on the performance gains due to the BD-RIS. Thus, the signal $r_{k}\in\mathbb{C}$ received at user $k$ can be expressed as
@@
where $\mathbf{r}\in\mathbb{C}^{K\times 1}$, $\mathbf{H}_{\mathrm{RX}}\in\mathbb{C}^{K\times R}$, $\mathbf{w}_{n}\in\mathbb{C}^{R\times 1}$ with $n\in\{1,2,\dots,N\}$, and $\mathbf{v}_{m}\in\mathbb{C}^{N\times 1}$ with $m\in\{1,2,\dots,K\}$, denote the received signal vector, BD-RIS-to-user $k$ channel matrix, BS-to-BD-RIS channel vector, and beamforming vector, respectively.
@@ II-A Scattering Matrix Definition
    Single-connected BD-RIS: Equivalent to the conventional D-RIS, where each RE is independently connected to a single grounded reconfigurable impedance with no interconnections among REs. Consequently, the resulting scattering matrix is inherently diagonal, which naturally satisfies the symmetry constraints, while also being unitary due to the lossless nature [^6] According to microwave network theory, a lossless reconfigurable impedance network must satisfy a unitary scattering matrix condition, thereby ensuring energy conservation [4]. Furthermore, the reciprocity property, which enforces symmetry of the BD-RIS scattering matrix, follows directly from electromagnetic reciprocity and microwave network theory. In particular, any passive, linear and time-invariant structure exhibits symmetric transmission characteristics, $i.e.$, the coupling between any two ports of the BD-RIS is identical in both directions, which is mathematically captured by the symmetric scattering matrix [37].. The aforementioned constraints are described as
@@
    Group-connected BD-RIS: In this intermediate architecture the REs are partitioned into $G$ disjoint groups, each consisting of $R_{G}=\frac{R}{G}$ elements. Within each group, all REs are interconnected, enabling waves impinging on one element to be reflected by all others in the same group. Apart from the intra-group connections, no other link exists. Consequently, the scattering matrix can be represented as a block-diagonal matrix by ordering its elements group-by-group, such that each block corresponds to one group and must independently satisfy the unitary and symmetry constraints. Notably, the group-connected architecture generalizes the previous two cases, reducing to the fully-connected case when $G=1$ and to the single-connected case when $G=R$. Denoting the scattering matrix of groups $g\in\{1,2,\ldots,G\}$ by $\mathbf{\Theta}_{g}\in\mathbb{C}^{R_{G}\times R_{G}}$, the overall structure of the group-connected architecture is expressed by
    $\mathbf{\Theta}=\mathrm{blkdiag}(\mathbf{\Theta}_{1},\ldots,\mathbf{\Theta}_{G}),$ (7)
    and the constraints on the group-connected BD-RIS scattering matrix are given by
    $\hskip-21.52771pt{\mathcal{S}_{\text{GC}_{1}}\!\!=\!\left\{\mathbf{\Theta}:\mathbf{\Theta}_{g}\!=\!\mathbf{\Theta}_{g}^{T},\forall g\right\}\!,\!\!\!}$ (8)
@@
    Here $\mathbf{\Theta}_{g}$ is formally defined as
    where $g\in\{1,2,\dots,G\}$. Figure 2, seen at the top of the next page, provides an illustrative example demonstrating the impact of beyond-diagonal connectivity on the scattering behavior of the BD-RIS.
(a) Single-Connected
(b) Group-Connected (group size 4)
(c) Fully-connected
Figure 2: Beampattern illustrating the responses of a BD-RIS with different levels of connectivity. The plots are shown in terms of the auxiliary quantities $u\triangleq\sin(\phi)\cos(\theta)$ and $v\triangleq\sin(\phi)\sin(\theta)$, where $\phi$ and $\theta$ are the azimuth and elevation angles, swept in the ranges $\pi,\pi]$ and $\pi/2,\pi/2]$, respectively. The blue lines correspond to the directions of two users, located at $(\phi_{1},\theta_{1})=(-40^{\circ},10^{\circ})$ and $(\phi_{2},\theta_{2})=(-20^{\circ},10^{\circ})$.